OpenStudy (anonymous):

how would i solve for x in the equation square root x-98+09=18

OpenStudy (cwrw238):

your question is not clear . What exaclty is the square root of ? Just the x or x - 98 or x - 98 + 9 ?

OpenStudy (anonymous):

you square root the whole equation

Nnesha (nnesha):

take square both side

OpenStudy (cwrw238):

also it looks like 9.8 and 0.9 not 98 and 9

OpenStudy (anonymous):

i know im sorry they are just marks on my paper. the numbers are the ones i told you earlier

OpenStudy (cwrw238):

well as Nnesha said if you square both sides the solution becomes straight forward

OpenStudy (cwrw238):

for example the square of sqrt( x + 3) is simply x + 3

OpenStudy (anonymous):

so would my first step be to +98 to both sides?

OpenStudy (cwrw238):

no you square sqrt(x + 98 + 09) and square 18

OpenStudy (cwrw238):

* that's sqrt(x - 98 + 09)

OpenStudy (cwrw238):

the square of the square root of something is the something

OpenStudy (anonymous):

so what would that equal out to? i know the square root of 18 would be 4.2

OpenStudy (cwrw238):

NO you square the 18

OpenStudy (cwrw238):

18*18

OpenStudy (anonymous):

so it would be 324

OpenStudy (cwrw238):

right and what is on the left hand side of the =?

OpenStudy (anonymous):

x-98+9

OpenStudy (cwrw238):

right so we have x - 98 + 9 = 324

OpenStudy (cwrw238):

-98 + 9 = -89 so x - 89 = 324 can you continue?

OpenStudy (anonymous):

so then you would add 89 to both sides right??

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

so x=413

OpenStudy (anonymous):

i have one more question? would it be extraneous

OpenStudy (cwrw238):

that correct answer sorry gotta go right now.